OpenStudy (jaydamarie1):

Solve for m. m + 2 = 17 8 15 19 34

Elsa213 (elsa213):

17-2=? 17-2=15 m=15

OpenStudy (jaydamarie1):

Solve for k. k−45=18k−45=18 −63−63 −27−27 27 63

OpenStudy (falconmaster52):

m = ± √17 = ± 4.1231 its what i got

OpenStudy (3mar):

so you just need the answer?